OpenStudy (firejay5):

For problem 39 - 41, use the following information: Adam and his family are planning to rent a midsize car for a one - day trip. In the Standard Rental Plan, they can rent a car or $52 per day + 23 cents per mile. In the Deluxe Rental Plan, the can rent a car for $80 per day with unlimited mileage. 40. Estimate the break - even point of the rental costs.

OpenStudy (saifoo.khan):

SRP: y = 0.23x + 52 DRP: y = 80

OpenStudy (firejay5):

ok

OpenStudy (firejay5):

@saifoo.khan What would 41 be?

OpenStudy (saifoo.khan):

y = 0.23x + 52 y = 0.23(15) + 52

OpenStudy (firejay5):

where did you get 15

OpenStudy (saifoo.khan):

150* sorry. typo

OpenStudy (firejay5):

y = 86.5

OpenStudy (saifoo.khan):

So it's 86.5 They would've chose Option B. since it's $80

OpenStudy (firejay5):

Is choice B cheaper or what

OpenStudy (saifoo.khan):

Ofcourse.

OpenStudy (firejay5):

sorry okay

OpenStudy (firejay5):

what's the break even point of the plans???

OpenStudy (saifoo.khan):

i think it's 80. not sure.

OpenStudy (saifoo.khan):

Or it can be: 0.23x + 52 =80

OpenStudy (firejay5):

will that get the break even point?

OpenStudy (saifoo.khan):

I'm not sure.. sorry.

OpenStudy (firejay5):

well thanks! @saifoo.khan